OpenStudy (anonymous):

help... krypton at 500 torr in a 250 ml container and helium at 950 torr in a 450 ml flask both at the same temperature are mixed by opening a stop wingspan in a narrow tube connecting the containers . calculate the total final pressure. a. 790 torr b. 670 torr

OpenStudy (anonymous):

you already are given the pressures of the two. P total = p1+p2

OpenStudy (anonymous):

Remember, if you mix the two containers, the volume changes too. Find the partial pressures and take the sum. So: (950torr x 450mL)/700mL = about 611 (500torr x 250mL)/700mL = about 179 611 + 179 = 790torr should be a.

OpenStudy (anonymous):

tnx a lot.. :)